OpenStudy (anonymous):

Simplify ((x+6)/12)/((x-8)/10)

OpenStudy (mimi_x3):

Is it: \[\huge \frac{\frac{\frac{x+6}{12}}{x-8}}{10} \]

OpenStudy (anonymous):

Yep

OpenStudy (mimi_x3):

rewrite it as: \[\large\frac{x+6}{12}*\frac{10}{x-8}\]

OpenStudy (mimi_x3):

Do you know why?

OpenStudy (anonymous):

No :/ plz explain :)

OpenStudy (mimi_x3):

hm..it's some recipocal rule http://www.themathpage.com/alg/reciprocals.htm

OpenStudy (mimi_x3):

i dont know how to explain sorry; haven't done these in a while i just know how to do it

OpenStudy (anonymous):

well, what do i do to simplify it after what u put above?

OpenStudy (mimi_x3):

well, \[\frac{10(x+6)}{12(x-8)}\]

OpenStudy (mimi_x3):

@lgbasallote: Might be able to help

OpenStudy (anonymous):

k

OpenStudy (lgbasallote):

i think @Mimi_x3 has done well

OpenStudy (mimi_x3):

\[\frac{10x+60}{12x-96}\]

OpenStudy (mimi_x3):

No; i don't know how to explain algebra lol

OpenStudy (anonymous):

yes thank you both

OpenStudy (mimi_x3):

It's not simplified yet

OpenStudy (mimi_x3):

wait.. \[\frac{10(x+6)}{12(x-8)}\] \[\frac{5(x+6)}{6(x-8)}\]